# MedQARo: A Large-Scale Benchmark for Medical Question Answering in Romanian

### v2

@@ 6 Conclusion
In this paper, we introduced MedQARo, the first large-scale benchmark for Romanian medical question answering. The new benchmark is the result of a meticulous data collection and annotation process carried out by experts in the medical domain, who constructed a dataset of 102,646 QA pairs. The QA samples are grounded in a collection of 1,011 medical case summaries of cancer patients. We conducted comprehensive experiments with several LLM families to assess the impact of various aspects, namely prompt structure and length, language vs. domain vs. general pretraining, and zero-shot prompting vs. LoRA-based fine-tuning. Our empirical results suggested some interesting findings. First of all, we observed that long prompts are not always helpful, degrading performance even if epicrises are typically long documents. Second of all, we found that generic pretraining can lead to better results, suggesting that prior language or domain adaption can actually produce LLMs with reduced generalization capabilities. Third of all, we determined that fine-tuning on our specific task significantly outperforms zero-shot prompting. Yet, even the best model was only able to correctly answer about $70\%$ of the test questions, indicating that MedQARo is a challenging benchmark.
@@
Acknowledgments
This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906.

### v3

@@ 6 Conclusion
In this paper, we introduced MedQARo, the first large-scale benchmark for Romanian medical question answering. The new benchmark is the result of a meticulous data collection and annotation process carried out by experts in the medical domain, who constructed a dataset of 105,880 QA pairs. The QA samples are grounded in a collection of 1,242 medical case summaries of cancer patients. We conducted comprehensive experiments with several LLM families to assess the impact of various aspects, namely prompt structure and length, language vs. domain vs. general pretraining, and zero-shot prompting vs. LoRA-based fine-tuning. Our empirical results suggested some interesting findings. First of all, we observed that long prompts are not always helpful, degrading performance even if epicrises are typically long documents. Second of all, we found that generic pretraining can lead to better results, suggesting that prior language or domain adaption can actually produce LLMs with reduced generalization capabilities. Third of all, we determined that fine-tuning on our specific task significantly outperforms zero-shot prompting. Yet, even the best model was only able to correctly answer about $70\%$ of the in-domain test questions and $40\%$ of the cross-domain test questions, indicating that MedQARo is a challenging benchmark.
@@
Declarations
Funding This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 351416. This work was also supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-0090, within PNCDI IV.
Conflict of interest The authors have no conflicts of interest to declare that are relevant to the content of this article.
Ethics approval The database was approved by the Ethics Committee of Colţea Clinical Hospital where the research was conducted, in accordance with Decision No. 19091, dated October 5th, 2021. The database was also approved by the Ethics Committee of “Dan Furtună” Medical Center, as per Decision No. 1323, dated 11th December, 2025.
Consent to participate The use of anonymized data for research purposes is based on the informed consent of patients, as recorded in the medical charts of Colţea Clinical Hospital and “Dan Furtună” Medical Center, and signed individually by each patient at admission. The entire process complies with the requirements of the General Data Protection Regulation (GDPR), the principles of the Declaration of Helsinki, and the standards of good medical and research practices.
Consent for publication The authors give their consent for publication.
Authors’ contributions A.C. Rogoz performed the experiments and wrote the initial draft. R.T. Ionescu designed the study and revised the manuscript. A.V. Anghel and I.L. Antone-Iordache participated in data collection and curation and wrote the initial draft. S. Coniac and A.I. Ionescu participated in data collection and curation and revised the manuscript.
Availability of data and materials The dataset has been made publicly available for non-commercial use at https://github.com/ana-rogoz/MedQARo.
Code availability The code has been made publicly available for non-commercial use at https://github.com/ana-rogoz/MedQARo.
Open Access This article is licensed under a Creative Commons Attribution 4.0 International License, which permits use, sharing, adaptation, distribution and reproduction in any medium or format, as long as you give appropriate credit to the original author(s) and the source, provide a link to the Creative Commons license, and indicate if changes were made. The images or other third party material in this article are included in the article’s Creative Commons license, unless indicated otherwise in a credit line to the material. If material is not included in the article’s Creative Commons license and your intended use is not permitted by statutory regulation or exceeds the permitted use, you will need to obtain permission directly from the copyright holder. To view a copy of this license, visit https://creativecommons.org/licenses/by/4.0/deed.en.